Program Information
Recent Log Entries
- APPROVED Wien filters high voltage evaluation
- COMMENT ON Wien filters high voltage evaluation
- COMMENT ON RF 1L08 Upgrade network infrastructure installation
- ced2Epics modified EPICS PVs
- iocsoftnl08rf restored
- COMMENT ON Hall A Dump Camera Box Shielding
- iocsoftRadNDXE401 restored
- COMMENT ON MO Outage in Preparation for Bldg 85 Panel PMs
- COMMENT ON SL 1427 MHz Driveline Amplifier Work
- COMMENT ON 26/27 SAM Power Supply Flushing
- COMMENT ON East Arc Magnet Flow Measurements
- iocsoftwatlowtcdev1 restored
- iocsoftwatlowtcdev1 restored
- COMMENT ON Plasma processing during the Fall 2026 SAM
- iocsoftnl09rf restored











